You are informed that a company produces Sulfuric acid by combining two variable factors, labor (L) and Sulphur (S), with one fixed input, capital stock (K0). The company pays $2 an hour for labor and $1 for Sulphur; the fixed costs associated with its capital stock total $150 a period.


(a) Write out the general form of the Company’s total costs


(b) Suppose the Company’s short-run production function is given by the following

Q = c (L, S, K0) = 2LP

Where each unit of Q represents a thousand products.

Determine the minimum total cost at which the company could produce 100,000 products.


(c) To achieve minimum cost, what quantities of labor and plastic would the company have to use?



1
Expert's answer
2021-06-16T04:29:22-0400

a)     The formula to calculate total cost is:

TC = TFC + TVC

Where,

TC =total cost

TFC =total fixed cost

TVC =total variable cost


b)     Given that.

Q = c (L, S, K0) = 2LP

Then,

100000/1000=2(2*P)

               100=4p

                   P=25

The minimum production cost will be 150+25= $175


c) To achieve minimum cost, the company have to adopt lean manufacturing principles that can reduce manufacturing costs by increasing labor productivity, cutting production through-put times, reducing inventories, and cutting errors and scrap by as much as half.